# Gauges and Accelerated Optimization over Smooth and/or Strongly Convex Sets

### v3

@@ 5.1 Application to $p$-norm Ellipsoid Feasibility Problems
Find $\text{Find }x\text{ s.t. }x\in\{x\mid\|A_{1}x-b_{1}\|_{p_{1}}\leq\tau_{1}\}\cap\{x\mid\|A_{2}x-b_{2}\|_{p_{2}}\leq\tau_{2}\}\ ,$ (28)
@@
Figure 2: The minimum accuracy of $\max_{i}\{\frac{1}{2}\gamma_{S_{i},e_{i}}^{2}(x_{k})\}-\max_{i}\{\frac{1}{2}\gamma_{S_{i},e_{i}}^{2}(x^{*})\}$ of finding the intersection point of two smooth/strongly convex $p$-norm ellipsoids. Note that all points below the black line at objective value one correspond to feasible solutions.
@@ 5.2 Application to Trust Region Optimization Problems
Finally, we compare the best of the previous first-order method (the accelerated generalized gradient method) against standard solvers for synthetic constrained minimization problems. We consider constrained quadratic minimization problems over a $p$-norm ellipsoidal constraint set
@@
For our projection-free method, we apply the accelerated generalized gradient method to the square of the radial reformulation (7) to benefit from any smoothness and/or strong convexity present in the constraints. (For this radial transformation to be well-defined, we need the origin to be feasible. To ensure this, we compute $e$ approximately solving $Ax=b$ via $30$ iteration of the conjugate gradient method and then translate the problem to place $e$ at the origin.) As in Section 5.1, we tune any algorithmic parameters (e.g., $\eta,L,\mu$) for each considered method to optimize final performance to two significant digits.
@@
where stepsizes tuned for each of the two cases of $\alpha_{k}=\begin{cases}\eta_{1}/\sqrt{k+10}&\text{if }g(x_{k})\leq 0\\ \eta_{2}/\sqrt{k+10}&\text{else}\end{cases}$. Second, we consider directly maximizing the problem’s Lagrangian $\mathcal{L}(x)=f(x)-\lambda^{*}g(x)$ assuming the dual optimal Lagrange multiplier $\lambda^{*}$ is given in advance. Then we compare with the following “ideal Lagrangian accelerated method”
@@
which has an equivalent unconstrained radially dual problem
@@
Note the second term above is exactly the gauge $\gamma_{\{x\mid\|Ax-b\|_{2}\leq 1\}}(y)$. As before, the smoothness and strong convexity of the objective and ellipsoid constraint corresponds to both terms above being smooth and strongly convex when squared. As a result, this can be solved at a linear rate by an accelerated projection-free method.
@@
Figure $4$Gurobi is limited to linear and quardratic programs, it is not able to run on $p=4$-norm ellipsoids. shows the first-order solvers have their performance fall off as the problem size grows. The second-order solver’s performance is matched by our radial method at all problem sizes, increasingly so as the problem dimension grows. This motivates future works, building more practical radial methods and testing their performance on real (non-synthetic) problem instances. Unlike the prior $p=2$ setting, the idealized Lagrangian accelerated method outperforms our radial method by a moderate factor (between two and five). Given this method’s benefited from knowledge of the optimal Lagrange multiplier, this is not unexpected.
@@ Acknowledgements
applying powers to vectors elementwise. Applying Lemma 6, $T_{p}$ then must be locally $\alpha(\bar{y})=\frac{p(p-1)}{\|\nabla f(\bar{y})\|_{2}\|\bar{y}-b\|_{\infty}^{2-p}}$-strongly convex if $p\leq 2$ and $\beta(\bar{y})=\frac{p(p-1)\|\bar{y}-b\|_{\infty}^{p-2}}{\|\nabla f(\bar{y})\|_{2}}$-smooth if $p\geq 2$. Our main theorems then give local smoothness and strong convexities for $\frac{1}{2}\gamma_{B_{p}}^{2}(y)=\frac{1}{2}\|y\|^{2}_{p}$. Computing the inf/sup of these quantities over all boundary points $\bar{y}$ gives the following strong convexity and smoothness constants
@@
For generic translated balls, here we provide relatively simple lower and upper bounds for $\mu$ and $L$ above. Namely, for strong convexity, above formula of $\mu$ can be lower bounded by
$\mu$ $\geq\inf_{\|x-b\|_{p}=1}\frac{\alpha(x)}{\frac{\nabla f(x)^{T}x}{\|\nabla f(x)\|_{2}}+\alpha(x)\|x\|^{2}_{2}}$
$\geq\inf_{\|x-b\|_{p}=1}\frac{1}{\|x\|_{2}+\|x\|^{2}_{2}}(p-1)n^{\frac{1}{2}-\frac{1}{p}}$
$\geq\frac{1}{(1+\|b\|_{2})(2+\|b\|_{2})}(p-1)n^{\frac{1}{2}-\frac{1}{p}}\ ,$
where the first inequality is by the same lower bound as (17), the second inequality is by Cauchy-Schwartz and that $\alpha(x)\leq(p-1)n^{\frac{1}{2}-\frac{1}{p}}\leq 1$, and the last inequality is using norm bounds $\|x\|_{2}\leq\|x-b\|_{2}+\|b\|_{2}\leq\|x-b\|_{p}+\|b\|_{2}=1+\|b\|_{2}$ Then by Lemma 15, $\frac{1}{2}\gamma_{E_{p}}^{2}$ is $\frac{\lambda_{\min}(A^{T}A)}{(1+\|b\|_{2})(2+\|b\|_{2})}(p-1)n^{\frac{1}{2}-\frac{1}{p}}$-strongly convex. For smoothness, above formula of $L$ can be upper bounded by
$L\leq$ $\sup_{\|x-b\|_{p}=1}\left(\frac{\frac{\nabla f(x)^{T}x}{\|\nabla f(x)\|_{2}}+\beta(x)\|x\|^{2}_{2}}{\left(\frac{\nabla f(x)^{T}x}{\|\nabla f(x)\|_{2}}\right)^{3}}\right)$
$=$ $\sup_{\|x-b\|_{p}=1}\left(\frac{p^{2}\nabla f(x)^{T}x\|x-b\|^{2(p-1)}_{2(p-1)}+p^{3}(p-1)\|x-b\|^{p-2}_{\infty}\|x\|^{2}_{2}\|x-b\|^{2(p-1)}_{2(p-1)}}{\left(\nabla f(x)^{T}x\right)^{3}}\right)$
$\leq$ $\frac{p^{2}(1-\|b\|_{p}^{p})+4p^{3}(p-1)n^{1-\frac{2}{p}}}{\left(1-\|b\|^{p}_{p}\right)^{3}}$
$\leq$ $\frac{9}{2}\left(\frac{p}{1-\|b\|_{p}^{p}}\right)^{3}(p-1)n^{1-\frac{2}{p}}\ ,$
where the first inequality is by the same upper bound as (18), the second inequality uses convexity as $\nabla f(x)^{T}x\geq f(x)-f(0)=1-\|b\|^{p}_{p}$ and the norm bounds $\|x-b\|_{\infty}\leq\|x-b\|_{2(p-1)}\leq\|x-b\|_{p}=1$, and $\|x\|_{2}\leq n^{\frac{1}{2}-\frac{1}{p}}\|x\|_{p}\leq n^{\frac{1}{2}-\frac{1}{p}}(\|x-b\|_{p}+\|b\|_{p})\leq 2n^{\frac{1}{2}-\frac{1}{p}}$, and the last inequality is using $1-\|b\|_{p}^{p}\leq\frac{1}{2}p(p-1)n^{1-\frac{2}{p}}$. Then by Lemma 15, $\frac{1}{2}\gamma_{E_{p}}^{2}$ is $\frac{9}{2}\lambda_{\max}(A^{T}A)\left(\frac{p}{1-\|b\|_{p}^{p}}\right)^{3}(p-1)n^{1-\frac{2}{p}}$-smooth.
@@ A.2.1 Proof of Proposition 1
The equivalence $\mathbf{(a)\Leftrightarrow(b)}$ is given by the equivalence in [17, Theorem 2.1.5 $\mathbf{(2.1.9)\Leftrightarrow(2.1.10)}$], and the equivalence $\mathbf{(a)\Leftrightarrow(c)}$ is given by the equivalence in [47, Proposition 12.60 $\mathbf{(e)\Leftrightarrow(d)}$]. The implication $\mathbf{(a)\Rightarrow(d)}$ is trivial as global smoothness implies local smoothness. The final implication $\mathbf{(d)\Rightarrow(b)}$ is nontrivial:
@@
If $f$ is globally $\Delta$-smooth and locally $\tilde{L}$-smooth everywhere, then $f$ is globally $\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)^{-1}$-smooth.
@@
Proof of Claim 23. We say $f$ has $L$-Lipschitz gradient at $x\in\mathcal{E}$ if for some $\eta>0$, every $y\in B(x,\eta)$ has $\|\nabla f(y)-\nabla f(x)\|\leq L\|y-x\|$. We say $f$ has pointwise $L$-Lipschitz gradient if it has $L$-Lipschitz gradient at every $x\in\mathcal{E}$. Note as a simple result, having pointwise $L$-Lipschitz gradient implies $f$ has uniformly $L$-Lipschitz gradient [48, Corollary 2.4]. Hence, our proof aims to show $f$ has pointwise $\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)^{-1}$-Lipschitz gradient. To that end, consider some $x\in\mathcal{E}$ and let $\eta>0$ be such that condition $\mathbf{(d)}$ holds at $x$.
@@
Therefore $\|\nabla f(y)-\nabla f(x)\|\leq\tilde{L}\eta$ since $(\nabla f(y)-\nabla f(x))^{T}(y-x)\leq\|\nabla f(y)-\nabla f(x)\|\eta/2$ by Cauchy-Schwartz. Thus we can consider $z=x+(\nabla f(y)-\nabla f(x))/\tilde{L}\in B(x,\eta)$. Then (37) implies
$f(y)\leq f(x)+\nabla f(y)^{T}(y-x)+\frac{1}{2\tilde{L}}\|\nabla f(y)-\nabla f(x)\|^{2}\ .$ (38)
Combining (36) and (38) gives $(\nabla f(y)-\nabla f(x))^{T}(y-x)\geq\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)\|\nabla f(y)-\nabla f(x)\|^{2}.$ Again applying Cauchy-Schwartz gives the desired Lipschitz gradient condition at $x$, from which we conclude $f$ has $\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}$-Lipschitz gradient globally.
@@ A.2.2 Proof of Proposition 2
The equivalence $\mathbf{(a)\Leftrightarrow(b)}$ is given by the equivalence in [49, Theorem 5.24 $\mathbf{(ii)\Leftrightarrow(iii)}$]. The equivalence $\mathbf{(a)\Leftrightarrow(c)}$ follows by combining a few standard results. Note $h$ being $\mu$-strongly convex is equivalent to the Fenchel conjugate $h^{*}$ being $\frac{1}{\mu}$-smooth by [47, Theorem 12.60 $\mathbf{(a)\Leftrightarrow(b)}$]. Moreover, $h^{*}$ is $\frac{1}{\mu}$-smooth if and only if there exists a closed convex function $g$ such that $g+h^{*}=\frac{1}{2\mu}\|\cdot\|^{2}$ by [50, Lemma 4 $\mathbf{[1]\Leftrightarrow[2]}$]. This statement about sums of conjugates is equivalent to the existence of $g^{*}$ with the following inf-convolution $g^{*}\ \square\ h=\frac{\mu}{2}\|\cdot\|^{2}$, by [47, Theorem 11.23(a)].
Finally, we show the equivalence with the local strong convexity condition $\mathbf{(d)}$. The implication $\mathbf{(a)\Rightarrow(d)}$ is trivial as global strong convexity implies local strong convexity. Again the local to global implication is nontrivial: We leverage our previous local to global result for smooth functions from Proposition 1(d). From this, it suffices to show that $f$ being locally $\tilde{\mu}<\mu$-strongly convex implies $f^{*}$ is locally $1/\tilde{\mu}$-smooth since one can then conclude $f^{*}$ is $1/\mu$-smooth and hence $f$ is $\mu$-strongly convex by [47, Theorem 12.60 $\mathbf{(a)\Leftrightarrow(b)}$].

### v4

@@ 5.1 Application to $p$-norm Ellipsoid Feasibility Problems
$\text{Find }x\text{ s.t. }x\in\{x\mid\|A_{1}x-b_{1}\|_{p_{1}}\leq\tau_{1}\}\cap\{x\mid\|A_{2}x-b_{2}\|_{p_{2}}\leq\tau_{2}\}\ ,$ (32)
@@
Figure 2: The minimum accuracy of $\max_{i}\{\frac{1}{2}\gamma_{S_{i},e_{i}}^{2}(x_{k})\}-\max_{i}\{\frac{1}{2}\gamma_{S_{i},e_{i}}^{2}(x^{*})\}$ of finding the intersection point of two smooth/strongly convex $p$-norm ellipsoids. Note that all points below the black line at objective value $1/2$ correspond to feasible solutions.
@@ 5.2 Application to Trust Region Optimization Problems
Finally, we compare the best of the previous first-order methods (the accelerated generalized gradient method) against standard solvers for synthetic constrained minimization problems. We consider constrained quadratic minimization problems over a $p$-norm ellipsoidal constraint set
@@
For our projection-free method, we apply the accelerated generalized gradient method to the square of the radial reformulation (7) to benefit from any smoothness and/or strong convexity present in the constraints. (For this radial transformation to be well-defined, we need the origin to be feasible. To ensure this, we compute $e$ by approximately solving $Ax=b$ via $30$ iteration of the conjugate gradient method and then translate the problem to place $e$ at the origin.) As in Section 5.1, we tune any algorithmic parameters (e.g., $\eta,L,\mu$) for each considered method to optimize final performance to two significant digits.
@@
where stepsizes are tuned for each of the two cases of $\alpha_{k}=\begin{cases}\eta_{1}/\sqrt{k+10}&\text{if }g(x_{k})\leq 0\\ \eta_{2}/\sqrt{k+10}&\text{else}\end{cases}$. Second, we consider directly maximizing the problem’s Lagrangian $\mathcal{L}(x)=f(x)-\lambda^{*}g(x)$ assuming the dual optimal Lagrange multiplier $\lambda^{*}$ is given in advance. Then we compare with the following “ideal Lagrangian accelerated method”
@@
which has an equivalent unconstrained radially dual problem  [^4] This is computed directly by definition, refer to [2] Section 2.1.
@@
Note the second term above is exactly the gauge $\gamma_{\{x\mid\|Ax-b\|_{2}\leq 1\}}(y)$. As before, the smoothness and strong convexity of the objective and ellipsoid constraint correspond to both terms above being smooth and strongly convex when squared. As a result, this can be solved at a linear rate by an accelerated projection-free method.
@@
Figure 4555Gurobi is limited to linear and quardratic programs, it is not able to run on $p=4$-norm ellipsoids. shows the first-order solvers have their performance fall off as the problem size grows. The second-order solver’s performance is matched by our radial method at all problem sizes, increasingly so as the problem dimension grows. This motivates future works, building more practical radial methods and testing their performance on real (non-synthetic) problem instances. Unlike the prior $p=2$ setting, the idealized Lagrangian accelerated method outperforms our radial method by a moderate factor (between two and five). Given this method’s benefited from knowledge of the optimal Lagrange multiplier, this is not unexpected.
@@ Acknowledgements
A useful lemma of the equivalence of $\ell_{p}$-norm in finite dimensions  [47, Proposition 6.12] is given below
Lemma 23.
For $x\in\mathbb{R}^{n}$, if $0<p<q\leq\infty$, then $\|x\|_{q}\leq\|x\|_{p}\leq n^{1/p-1/q}\|x\|_{q}$.
@@
applying powers to vectors elementwise. Applying Lemma 6, $T_{p}$ then must be locally $\alpha(\bar{y})$-strongly convex and $\beta(\bar{y})$-smooth such that
$\text{for }p\leq 2\quad\alpha(\bar{y})=\frac{p(p-1)}{\|\nabla f(\bar{y})\|_{2}\,\|\bar{y}-b\|_{\infty}^{\,2-p}},\qquad\text{while for }p\geq 2\quad\beta(\bar{y})=\frac{p(p-1)\|\bar{y}-b\|_{\infty}^{\,p-2}}{\|\nabla f(\bar{y})\|_{2}}.$
Using $\|\bar{y}-b\|_{\infty}\leq 1$ and Lemma 23 gives the pointwise bounds
$\alpha(\bar{y})\ \geq\ \alpha_{*}:=(p-1)\,n^{\frac{1}{2}-\frac{1}{p}},\qquad\beta(\bar{y})\ \leq\ \beta^{*}:=(p-1)\,n^{\frac{1}{2}-\frac{1}{p}}.$
Our main theorems then give local smoothness and strong convexities for $\frac{1}{2}\gamma_{B_{p}}^{2}(y)=\frac{1}{2}\|y\|^{2}_{p}$. Computing the inf/sup of these quantities over all boundary points $\bar{y}$ gives the following strong convexity and smoothness constants
@@
For generic translated balls, here we provide relatively simple lower and upper bounds for $\mu$ and $L$ above.
Strong convexity. By the gauge strong convexity bound (the same inequality as in (18)),
$\mu\ \geq\ \inf_{\|x-b\|_{p}=1}\ \frac{\alpha(x)}{\frac{\nabla f(x)^{T}x}{\|\nabla f(x)\|_{2}}+\alpha(x)\|x\|_{2}^{2}}.$
Using $\frac{\nabla f(x)^{T}x}{\|\nabla f(x)\|_{2}}\leq\|x\|_{2}$ (Cauchy–Schwarz) and the monotonicity of $\alpha\mapsto\frac{\alpha}{A+\alpha t}$ for $A,t\geq 0$, we obtain
$\mu\ \geq\ \inf_{\|x-b\|_{p}=1}\ \frac{\alpha_{*}}{\|x\|_{2}+\alpha_{*}\,\|x\|_{2}^{2}}.$
Finally, $\|x\|_{2}\leq\|x-b\|_{2}+\|b\|_{2}\leq\|x-b\|_{p}+\|b\|_{2}=1+\|b\|_{2}$ yields the explicit bound
$\quad\mu\ \geq\ \frac{\alpha_{*}}{(1+\|b\|_{2})\ +\ \alpha_{*}(1+\|b\|_{2})^{2}}\ =\ \frac{(p-1)\,n^{\frac{1}{2}-\frac{1}{p}}}{(1+\|b\|_{2})\ +\ (p-1)\,n^{\frac{1}{2}-\frac{1}{p}}(1+\|b\|_{2})^{2}}\ .\quad$ (40)
By Lemma 15, for $E_{p}=\{y:\ Ay-b\in B_{p}\}$ we get
$\text{$\frac{1}{2}\gamma_{E_{p}}^{2}$ is }\quad\lambda_{\min}(A^{T}A)\times\mu\text{-strongly convex,}$
so (40) multiplied by $\lambda_{\min}(A^{T}A)$ provides a global constant.
Smoothness. From the gauge smoothness bound (the same inequality as in (20)), on $\partial T_{p}$,
[…]
Using $\|x-b\|_{\infty}\leq 1$, $\|x-b\|_{2(p-1)}\leq 1$, Lemma 23, and convexity $\nabla f(x)^{T}x\geq f(x)-f(0)=1-\|b\|_{p}^{p}$, we get
$\quad L\ \leq\ \frac{p^{2}\,(1-\|b\|_{p}^{p})\ +\ p^{3}(p-1)\,n^{\,1-\frac{2}{p}}\,\big(1+\|b\|_{p}\big)^{2}}{(1-\|b\|_{p}^{p})^{3}}\ .\quad$ (41)
If we only want a simpler expression under the standing assumption $\|b\|_{p}<1$, then $(1+\|b\|_{p})^{2}\leq 4$ and $1-\|b\|_{p}^{p}\leq 1\leq\frac{1}{2}p(p-1)n^{1-\frac{2}{p}}$, hence (41) reduces to
[…]
Again, Lemma 15 multiplies this by $\lambda_{\max}(A^{T}A)$ for the set $E_{p}$.
@@ A.2.1 Proof of Proposition 1
The equivalence $\mathbf{(a)\Leftrightarrow(b)}$ is given by the equivalence in [17, Theorem 2.1.5 $\mathbf{(2.1.9)\Leftrightarrow(2.1.11)}$], and the equivalence $\mathbf{(a)\Leftrightarrow(c)}$ is given by the equivalence in [48, Proposition 12.60 $\mathbf{(e)\Leftrightarrow(d)}$]. The implication $\mathbf{(a)\Rightarrow(d)}$ is trivial as global smoothness implies local smoothness. The final implication $\mathbf{(d)\Rightarrow(b)}$ is nontrivial:
@@
If $f$ is globally $\Delta$-smooth and locally $\tilde{L}-$smooth everywhere, then $f$ is globally $\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)^{-1}$-smooth. [^6] Both $\Delta$ and $\tilde{L}$ can be $\infty$.
@@
Proof of Claim 24. We say $f$ has $L$-Lipschitz gradient at $x\in\mathcal{E}$ if for some $\eta>0$, every $y\in B(x,\eta)$ has $\|\nabla f(y)-\nabla f(x)\|\leq L\|y-x\|$. We say $f$ has pointwise $L$-Lipschitz gradient if it has $L$-Lipschitz gradient at every $x\in\mathcal{E}$. Note as a simple result, having pointwise $L$-Lipschitz gradient implies $f$ has uniformly $L$-Lipschitz gradient [49, Corollary 2.4]. Hence, our proof aims to show $f$ has pointwise $\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)^{-1}$-Lipschitz gradient. To that end, consider some $x\in\mathcal{E}$ and let $\eta>0$ be such that condition $\mathbf{(d)}$ holds at $x$.
@@
Therefore $\|\nabla f(y)-\nabla f(x)\|\leq\tilde{L}\eta$ since $(\nabla f(y)-\nabla f(x))^{T}(y-x)\leq\|\nabla f(y)-\nabla f(x)\|\eta/2$ by Cauchy-Schwarz. Thus we can consider $z=x+(\nabla f(y)-\nabla f(x))/\tilde{L}\in B(x,\eta)$. Then (43) implies
$f(y)\leq f(x)+\nabla f(y)^{T}(y-x)-\frac{1}{2\tilde{L}}\|\nabla f(y)-\nabla f(x)\|^{2}\ .$ (44)
Combining (42) and (44) gives $(\nabla f(y)-\nabla f(x))^{T}(y-x)\geq\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)\|\nabla f(y)-\nabla f(x)\|^{2}.$ Again applying Cauchy-Schwarz gives the desired Lipschitz gradient condition at $x$, from which we conclude $f$ has $\left(\frac{1}{2\Delta}+\frac{1}{2\tilde{L}}\right)^{-1}$-Lipschitz gradient globally.
@@ A.2.2 Proof of Proposition 2
The equivalence $\mathbf{(a)\Leftrightarrow(b)}$ is given by the equivalence in [50, Theorem 5.24 $\mathbf{(ii)\Leftrightarrow(iii)}$]. The equivalence $\mathbf{(a)\Leftrightarrow(c)}$ follows by combining a few standard results. Note $h$ being $\mu$-strongly convex is equivalent to the Fenchel conjugate $h^{*}$ being $\frac{1}{\mu}$-smooth by [48, Theorem 12.60 $\mathbf{(a)\Leftrightarrow(b)}$]. Moreover, $h^{*}$ is $\frac{1}{\mu}$-smooth if and only if there exists a closed convex function $g$ such that $g+h^{*}=\frac{1}{2\mu}\|\cdot\|^{2}$ by [51, Lemma 4 $\mathbf{[1]\Leftrightarrow[2]}$]. This statement about sums of conjugates is equivalent to the existence of $g^{*}$ with the following inf-convolution $g^{*}\ \square\ h=\frac{\mu}{2}\|\cdot\|^{2}$, by [48, Theorem 11.23(a)].
Finally, we show the equivalence with the local strong convexity condition $\mathbf{(d)}$. The implication $\mathbf{(a)\Rightarrow(d)}$ is trivial as global strong convexity implies local strong convexity. Again the local to global implication is nontrivial: We leverage our previous local to global result for smooth functions from Proposition 1(d). From this, it suffices to show that $f$ being locally $\tilde{\mu}<\mu$-strongly convex implies $f^{*}$ is locally $1/\tilde{\mu}-$smooth since one can then conclude $f^{*}$ is $1/\mu$-smooth and hence $f$ is $\mu$-strongly convex by [48, Theorem 12.60 $\mathbf{(a)\Leftrightarrow(b)}$].